# Intrinsic Mixed-state Quantum Topological Order

### v2

@@ IV.1 Decohered Kitaev honeycomb model
In last section we provide a general route to generalize the construction in Section II to obtain other intrinsic mixed-state quantum TO. In this section we give two more examples as applications. We first discuss the generalization to the Kitaev honeycomb model: $H=-J_{x}\sum_{x\text{-bonds }}\sigma_{j}^{x}\sigma_{k}^{x}-J_{y}\sum_{y\text{-bonds }}\sigma_{j}^{y}\sigma_{k}^{y}-J_{z}\sum_{z\text{-bonds }}\sigma_{j}^{z}\sigma_{k}^{z}-\sum_{i}\vec{h}\cdot\vec{\sigma}_{i}$ (with $|\vec{h}|\ll|J_{\mu}|,\mu=x,y,z$). This model can be exactly solved by mapping it to Majorana fermions coupled to static $\mathbb{Z}_{2}$ gauge fields Kitaev (2003). It is shown that the ground state $\rho_{0}$ of this model can realize Abelian $\mathbb{Z}_{2}$ TO, non-Abelian Ising TO, as well as a gapless $\mathbb{Z}_{2}$ spin liquid phase. All the three phases have deconfined fermion excitations.
@@
To obtain intrinsic mixed-state quantum TO, we construct the following channel:
@@
This channel leads to decoherence of the fermions but preserves the $\mathbb{Z}_{2}$ gauge flux. In other words, it preserves the anomalous $1$-form symmetry, with generators $W^{f}_{p}=\prod_{\langle ij\rangle\in p}\sigma^{z}_{1}\sigma^{y}_{2}\sigma^{x}_{3}\sigma^{z}_{4}\sigma^{y}_{5}\sigma^{x}_{6}$ Liu et al. (2024). Therefore, the resulting mixed states must be LRE, and support deconfined fermionic excitations. In the maximally decohered case $p_{f}=\frac{1}{2}$, the $f$ particles are heated to infinite temperature, so all of the three ground-state phases will end up in the maximally mixed state in the zero-flux sector ($W^{f}_{p}=1,\forall p$), which belongs to the same intrinsic mixed-state quantum TO as that constructed in Section II. Actually, at $p_{f}=\frac{1}{2}$, this state can be obtained by applying a Hadamart gate on all vertical links to (11). Therefore, the decohered honeycomb model constructed here is also characterized by a nonzero TEN. We note that a similar model in the context of Lindblad equations was constructed in Hwang (2023), but the LRE nature was not uncovered.
@@ IV.2 Decohered double semion model
As another example, we construct intrinsic mixed-state quantum TO from the double semion TO. The anyon content of the double semion TO is $\mathcal{A}=\{1,s,\bar{s},s\bar{s}\}=\{1,s\}\times\{1,\bar{s}\}$, where $s$ is a semion, $\theta(s)=i$; $\bar{s}$ is an antisemion, $\theta(\bar{s})=-i$; and $s\bar{s}$ is a boson, $\theta(s\bar{s})=1$. The fusion rules are $s\times s=1,\bar{s}\times\bar{s}=1,s\times\bar{s}=s\bar{s}$ Levin and Wen (2005). Following the general strategy in Section III, we can proliferate the semion $s$ using noisy channels, and due to the trivial braiding between $s$ and $\bar{s}$, the $1$-form symmetry generated by $\bar{s}$ is preserved. Thus $\bar{s}$ remains a strongly deconfined excitation with well-defined anti-semionic statistics.
@@
$\mathcal{N}^{[em]}=\prod_{i}\mathcal{N}^{[em]}_{i},\ \mathcal{N}^{[em]}_{i}[\cdot]\equiv\sum_{n=0,1,2,3}p_{n}K^{n}_{i}\cdot K^{n\dagger}_{i}$ (29)
The Kraus operators $K_{i}$, $K^{\dagger}_{i}$ are the shortest string operators of $[em]$:
$K_{i}=\left\{\begin{aligned} &Z_{i}X_{i}\text{ for vertical link }i\\ &Z_{i}X^{-1}_{i}\text{ for horizontal link }i\end{aligned}\right.$ (30)
For simplicity, we take the initial state $\rho_{0}$ to be the maximally mixed state in the ground-state subspace of ${H}_{\text{DS}}$ and directly consider maximally decohered case $p_{0}=p_{1}=p_{2}=p_{3}=\frac{1}{4}$. Based on the analysis at the beginning of this section, it seems only the anti-semion $[e^{3}m]$ will remain strongly deconfined, which leads to a chiral anti-semion theory $\{1,[e^{3}m]\}$. However, the actual situation turns out to be even more intriguing.
First, we note that the $1$-form symmetry generated by $e^{3}m$ is indeed preserved because $[W^{e^{3}m}_{\gamma},K_{i}]=0$, for any closed $e^{3}m$ strings, which implies that $e^{3}m$ is a deconfined anyon. Examples of a shortest closed $e^{3}m$ string as well as an open $e^{3}m$ string is given in Fig.6(c). On the other hand, the $C_{i}$ terms in $H_{\text{DS}}$ do not commute with the Kraus operators. As a result, for any open $e^{2}m^{2}$ string $W^{e^{2}m^{2}}_{C}=\prod_{i\in C}C_{i}$, $\text{tr}(\rho W_{C}^{e^{2}m^{2}})$ becomes 00 for $\rho=\mathcal{N}^{[em]}[\rho_{0}]$. This means that the $e^{2}m^{2}$ anyons are revived from the Bose-Einstein condensate and become a detectable anyon. Notably, the $e^{2}m^{2}$ still proliferate classically, which is very different from Bose-Einstein condensation as we noted previously. Moreover, due to the strong $1$-form symmetry generated by $e^{3}m$ and the fusion rule $e^{3}m\times e^{3}m=e^{2}m^{2}$, $e^{2}m^{2}$ must become a deconfined anyon. Therefore, the remaining strongly deconfined anyons in $\rho$ form a $\mathbb{Z}_{4}$ group $\{1,e^{3}m,e^{2}m^{2},em^{3}\}$. We note that $e^{3}m$ and $em^{3}$ are both anti-semions and $e^{2}m^{2}$ is a transparent boson, meaning that it has trivial braiding with all strongly deconfined anyons. Anyon theories with transparent anyons are known as non-modular anyon theories Kitaev (2006); Ellison et al. (2023). It is widely believed that non-modular anyon theories cannot be realized by local gapped Hamiltonians in 2D bosonic systems Levin (2013); Kong and Wen (2014). This implies the lack of a pure-state counterpart of the mixed-state TO; thus it is indeed intrinsically mixed. Notably, the intrinsic mixed-state quantum TO constructed in Section II is also non-modular, with strongly deconfined anyons $\{1,f\}$. One crucial difference is that here the intrinsic mixed-state quantum TO does have a quantum memory. The undamaged part of the stored information is manipulated by the logical operators shown in Fig.7.
@@
$W^{e^{3}m}_{\tilde{\gamma_{x}}}W^{e^{3}m}_{\tilde{\gamma_{y}}}=-W^{e^{3}m}_{\tilde{\gamma_{y}}}W^{e^{3}m}_{\tilde{\gamma_{x}}}$ as a consequence of the nontrivial self-braiding statistics of $e^{3}m$. Besides, both $(W^{e^{3}m}_{\tilde{\gamma_{x}}})^{2},(W^{e^{3}m}_{\tilde{\gamma_{x}}})^{2}$ are elements of the stabilizer group defined by $H_{\text{DS}}$, thus acting trivially in the code space. Therefore, the intrinsic mixed-state quantum TO supports quantum memory with one and only one logical qubit.
@@
Remarkably, the anyon theory here is identical to the one obtained by incoherently proliferating the $em,e^{2}m^{2},e^{3}m^{3}$ in the $\mathbb{Z}_{4}$ toric code via the same channel $\mathcal{N}^{[em]}$ Sohal and Prem (2024); Ellison and Cheng (2024). The relation between $\mathbb{Z}_{4}$ toric code, double semion, and the non-modular anyon theory is summarized in Fig.8. However, it does not imply the final mixed states in the two models to be identical. Actually, the decohered $\mathbb{Z}_{4}$ toric code model (under $\mathcal{N}^{[em]}$) is fully characterized by the stabilizer group $G_{e^{3}m}=\langle\{W^{e^{3}m}_{p}\}\rangle$, which defines the code space $\mathcal{H}_{C}$:
@@
In the maximally decohered case, the final state is the maximally mixed state in $\mathcal{H}_{C}$. For the decohered double semion model, however, there are additional non-local stabilizers formed by products of $C_{i}$. Two such non-local stabilizers are depicted in Fig.7, and others non-local stabilizers can be obtained from these two via translation along the horizontal direction.
We summarize several surprising features of intrinsic mixed-state quantum TO revealed by this example. First, novel non-modular TO beyond the usual unitary modular tensor category description of 2+1D TO can be easily realized by Pauli stabilizer models under decoherence Sohal and Prem (2024); Ellison and Cheng (2024). Secondly, decoherence can sometimes give rise to new types of deconfined anyons that is absent in the anyon theory supported by the ground-state TO. It further implies that some features of the mixed-state TO can go beyond the prediction based on the anyon data, including the field-theoretic description Lee et al. (2023); Bao et al. (2023).
Figure 8: Relations between $\mathbb{Z}_{4}$ toric code, double semion, and the non-modular anyon theory supported by intrinsic mixed-state quantum TO.
@@ V Discussion
Our work introduces a promising mechanism for creating novel topologically ordered phases in mixed states. We give two complementary perspectives to demonstrate such possibility. The first perspective is to look at what anyons are proliferated. One of our key observations is that while the ways of anyon condensation are limited for pure states, anyon proliferation in mixed states can occur in more general ways, offering new possibilities for topological order. In the three models studied in this work, we propose new types of topological order arising from incoherent proliferation of fermionic or semionic anyons in ground-state topological order, which drives an unconventional phase transition that does not resemble any anyon condensation transition in pure states. The other perspective is to look at what remains. We find that the existence of anomalous strong $1$-form symmetries can be viewed as a guiding principle to novel mixed-state TO. We give a general proof that anomalous strong 1-form symmetries implies the LRE nature of the mixed states, which manifests as deconfined anyons with nontrivial statistics. By analyzing the fusion rules and statistics of deconfined anyons, we show the possibility of realizing non-modular quantum TO in mixed states.
As is clear from the three examples, intrinsic mixed-state quantum TO can have or not have quantum memory, depending on whether the remaining deconfined anyons have nontrivial braiding statistics or only fermionic statistics Bao et al. (2023).
@@
We end with some open directions. Firstly, as already mentioned in Section III, it remains unclear whether $1$-form symmetries generated by fermions also guarantee bipartite long-range entanglement, or there exist counterexamples which are bipartite separable but have multipartite long-range entanglement Lessa et al. (2024). Secondly, a systematic classification of the intrinsic mixed-state quantum TO is still lacking, with two main difficulties. The first is how to treat the weakly deconfined anyons with the peculiar one-way braiding statistics. The other is how to generalize the discussion of anyon statistics to mixed states away from the fixed-point models, like the one discussed in II.4. Thirdly, it will be interesting to generalize our construction to get non-Abelian intrinsic mixed-state quantum TO, of which a simple route is to proliferate a subset of Abelian anyons in a known non-Abelian TO. The more intriguing case is to proliferate the non-Abelian anyons, in which case the decoherence-induced transition can be much harder to analyze. Finally, in our construction, we start from a topologically ordered state and get its descendants via noisy channels. It is also tempting to find systematic ways to prepare mixed-state topological order from short-range entangled mixed states, for example, by measurng mixed-state symmetry protected topological order Ma and Wang (2023); Lee et al. (2022b); Zhang et al. (2022); Ma et al. (2023); Lu et al. (2023).
@@ Acknowledgements.
Acknowledgments.—We are especially grateful to Ruihua Fan for many pieces of valuable advice and feedback on the manuscript. We also thank Tsung-Cheng Lu for pointing out to us the dependence of TEN on the choice of entanglement cut. We thank Yingfei Gu, Meng Cheng, Yu-An Chen, Zhen Bi, Jing-Yuan Chen, and He-Ran Wang for helpful discussions. This work is supported by NSFC under Grant No. 12125405. Zhengzhi Wu acknowledges the support of the Shuimu Fellow Foundation at Tsinghua University.

### v3

@@ IV.1 Decohered Kitaev honeycomb model
In the last section, we provide a general route to generalize the construction in Section II to obtain other intrinsic mixed-state TO. In this section, we give two more examples as applications. We first discuss the generalization to the Kitaev honeycomb model: $H=-J_{x}\sum_{x\text{-bonds }}\sigma_{j}^{x}\sigma_{k}^{x}-J_{y}\sum_{y\text{-bonds }}\sigma_{j}^{y}\sigma_{k}^{y}-J_{z}\sum_{z\text{-bonds }}\sigma_{j}^{z}\sigma_{k}^{z}-\sum_{i}\vec{h}\cdot\vec{\sigma}_{i}$ (with $|\vec{h}|\ll|J_{\mu}|,\mu=x,y,z$). This model can be exactly solved by mapping it to Majorana fermions coupled to static $\mathbb{Z}_{2}$ gauge fields Kitaev (2003). It is shown that the ground state $\rho_{0}$ of this model can realize Abelian $\mathbb{Z}_{2}$ TO, non-Abelian Ising TO, as well as a gapless $\mathbb{Z}_{2}$ spin liquid phase. All three phases have deconfined fermion excitations.
@@
To obtain intrinsic mixed-state TO, we construct the following channel:
@@
This channel leads to decoherence of the fermions but preserves the $\mathbb{Z}_{2}$ gauge flux. In other words, it preserves the anomalous $1$-form symmetry, with generators $W^{f}_{p}=\prod_{\langle ij\rangle\in p}\sigma^{z}_{1}\sigma^{y}_{2}\sigma^{x}_{3}\sigma^{z}_{4}\sigma^{y}_{5}\sigma^{x}_{6}$ Liu et al. (2024). Therefore, the resulting mixed states must be LRE, and support deconfined fermionic excitations. In the maximally decohered case $p_{f}=\frac{1}{2}$, the $f$ particles are heated to infinite temperature, so all of the three ground-state phases will end up in the maximally mixed state in the zero-flux sector ($W^{f}_{p}=1,\forall p$), which belongs to the same intrinsic mixed-state TO as that constructed in Section II. Actually, at $p_{f}=\frac{1}{2}$, this state can be obtained by applying a Hadamard gate on all vertical links to (11). Therefore, the decohered honeycomb model constructed here is also characterized by a nonzero TEN. We note that a similar model in the context of Lindblad equations was constructed in Hwang (2023), but the LRE nature was not uncovered.
@@ IV.2 Decohered double semion model
As another example, we construct intrinsic mixed-state TO from the double semion TO. The anyon content of the double semion TO is $\mathcal{A}=\{1,s,\bar{s},s\bar{s}\}=\{1,s\}\times\{1,\bar{s}\}$, where $s$ is a semion, $\theta(s)=i$; $\bar{s}$ is an antisemion, $\theta(\bar{s})=-i$; and $s\bar{s}$ is a boson, $\theta(s\bar{s})=1$. The fusion rules are $s\times s=1,\bar{s}\times\bar{s}=1,s\times\bar{s}=s\bar{s}$ Levin and Wen (2005). Following the general strategy in Section III, we can proliferate the semion $s$ using noisy channels, and due to the trivial braiding between $s$ and $\bar{s}$, the $1$-form symmetry generated by $\bar{s}$ is preserved. Thus $\bar{s}$ remains a strongly deconfined excitation with well-defined anti-semionic statistics.
@@
$\mathcal{N}^{[em]}=\prod_{i}\mathcal{N}^{[em]}_{i},\ \mathcal{N}^{[em]}_{i}[\cdot]\equiv\sum_{n=0,1,2,3}p_{n}K^{n}_{i}\cdot K^{n\dagger}_{i}.$ (29)
The Kraus operators $K_{i}$, $K^{\dagger}_{i}$ are the shortest string operators of $em$ anyons in the $Z_{4}$ toric code:
$K_{i}=\left\{\begin{aligned} &Z_{i}X_{i+\bm{\delta}}\text{ for vertical link }i\\ &Z_{i}X^{-1}_{i+\bm{\delta}}\text{ for horizontal link }i\end{aligned}\right.$ (30)
For simplicity, we take the initial state $\rho_{0}$ to be the maximally mixed state in the ground-state subspace of ${H}_{\text{DS}}$ and directly consider the maximally decohered case $p_{0}=p_{1}=p_{2}=p_{3}=\frac{1}{4}$. Based on the analysis at the beginning of this section, it seems only the anti-semion $[e^{3}m]$ will remain strongly deconfined, which leads to a chiral anti-semion theory $\{1,[e^{3}m]\}$. However, the actual situation turns out to be even more intriguing.
First, we note that the $1$-form symmetry generated by $e^{3}m$ is indeed preserved because $[W^{e^{3}m}_{\gamma},K_{i}]=0$, for any closed $e^{3}m$ strings, which implies that $e^{3}m$ is a deconfined anyon. Examples of a shortest closed $e^{3}m$ string as well as an open $e^{3}m$ string is given in Fig. 6(c). On the other hand, the $C_{i}$ terms in $H_{\text{DS}}$ do not commute with the Kraus operators. As a result, for any open $e^{2}m^{2}$ string $W^{e^{2}m^{2}}_{C}=\prod_{i\in C}C_{i}$, $\text{tr}(\rho W_{C}^{e^{2}m^{2}})$ becomes 00 for $\rho=\mathcal{N}^{[em]}[\rho_{0}]$. This means that the $e^{2}m^{2}$ anyons are revived from the Bose-Einstein condensate and become a detectable anyon. Notably, the $e^{2}m^{2}$ still proliferate classically, which is very different from Bose-Einstein condensation as we noted previously. Moreover, due to the strong $1$-form symmetry generated by $e^{3}m$ and the fusion rule $e^{3}m\times e^{3}m=e^{2}m^{2}$, $e^{2}m^{2}$ must become a deconfined anyon. Therefore, the remaining strongly deconfined anyons in $\rho$ form a $\mathbb{Z}_{4}$ group $\{1,e^{3}m,e^{2}m^{2},em^{3}\}$. We note that $e^{3}m$ and $em^{3}$ are both anti-semions and $e^{2}m^{2}$ is a transparent boson, meaning that its presence cannot be remotely detected via an Aharonov-Bohm measurement, i.e., a full braid of any strongly deconfined anyon around it only results in a unity phase factor. Anyon theories with transparent bosons/fermions are known as non-modular anyon theories Kitaev (2006); Ellison et al. (2023). It is widely believed that non-modular anyon theories cannot be realized by local gapped Hamiltonians in 2D bosonic systems Levin (2013); Kong and Wen (2014). This implies the lack of a pure-state counterpart of the mixed-state TO; thus it is indeed intrinsically mixed. Notably, the intrinsic mixed-state TO constructed in Section II is also non-modular, with strongly deconfined anyons $\{1,f\}$. One crucial difference is that here the intrinsic mixed-state TO does have a quantum memory. The undamaged part of the stored information is manipulated by the logical operators shown in Fig. 7.
@@
$W^{e^{3}m}_{\tilde{\gamma_{x}}}W^{e^{3}m}_{\tilde{\gamma_{y}}}=-W^{e^{3}m}_{\tilde{\gamma_{y}}}W^{e^{3}m}_{\tilde{\gamma_{x}}}$ as a consequence of the nontrivial self-braiding statistics of $e^{3}m$. Besides, both $(W^{e^{3}m}_{\tilde{\gamma_{x}}})^{2},(W^{e^{3}m}_{\tilde{\gamma_{x}}})^{2}$ are elements of the stabilizer group defined by $H_{\text{DS}}$, thus acting trivially in the code space. Therefore, the intrinsic mixed-state TO supports quantum memory with one and only one logical qubit.
@@
Remarkably, the anyon theory here is identical to the one obtained by incoherently proliferating the $em,e^{2}m^{2},e^{3}m^{3}$ in the $\mathbb{Z}_{4}$ toric code via the same channel $\mathcal{N}^{[em]}$ Sohal and Prem (2024); Ellison and Cheng (2024). The relation between $\mathbb{Z}_{4}$ toric code, double semion, and the non-modular anyon theory is summarized in Fig. 8. However, it does not imply that the final mixed states in the two models are identical. Actually, the decohered $\mathbb{Z}_{4}$ toric code model (under $\mathcal{N}^{[em]}$) is fully characterized by the stabilizer group $G_{e^{3}m}=\langle\{W^{e^{3}m}_{p}\}\rangle$, which defines the code space $\mathcal{H}_{C}$:
@@
In the maximally decohered case, the final state is the maximally mixed state in $\mathcal{H}_{C}$. For the decohered double semion model, however, there are additional non-local stabilizers formed by products of $C_{i}$. Two such non-local stabilizers are depicted in Fig. 7, and other non-local stabilizers can be obtained from these two via translation along the horizontal direction.
We summarize several surprising features of intrinsic mixed-state TO revealed by this example. First, novel non-modular TO beyond the usual unitary modular tensor category description of 2+1D TO can be easily realized by Pauli stabilizer models under decoherence Sohal and Prem (2024); Ellison and Cheng (2024). Secondly, decoherence can sometimes give rise to new types of deconfined anyons that is absent in the anyon theory supported by the ground-state TO. It further implies that some features of the mixed-state TO can go beyond the prediction based on the IR theory/anyon data of the original topological order (i.e., before decoherence), including the field-theoretic description Lee et al. (2023); Bao et al. (2023). Indeed, if we start from other lattice realizations of double semion topological order, e.g., the $\mathbb{Z}_{4}$ Pauli stabilizer code introduced in Ellison et al. (2023) (which is a variant of (27)) or the twisted $\mathbb{Z}_{2}$ lattice gauge theory in Dijkgraaf and Witten (1990); Levin and Gu (2012); Hu et al. (2013); Magdalena de la Fuente et al. (2021); Song et al. (2024), and incoherently proliferate semions via quantum channels, we would get a chiral anti-semion theory $\{1,\bar{s}\}$ instead. The corresponding channels can be constructed by just taking the Kraus operators to be the semion string operators. For the explicit expression of the shortest string operators of twisted $\mathbb{Z}_{2}$ gauge theory, we refer the readers to the supplemental material of Song et al. (2024).
Figure 8: Relations between $\mathbb{Z}_{4}$ toric code, double semion, and the non-modular anyon theory supported by intrinsic mixed-state TO.
@@ V Discussion
Our work introduces a promising mechanism for creating novel topologically ordered phases in mixed states. We give two complementary perspectives to demonstrate such a possibility. The first perspective is to look at what anyons are proliferated. One of our key observations is that while the ways of anyon condensation are limited for pure states, anyon proliferation in mixed states can occur in more general ways, offering new possibilities for topological order. In the three models studied in this work, we propose new types of topological order arising from incoherent proliferation of fermionic or semionic anyons in ground-state topological order, which drives an unconventional phase transition that does not resemble any anyon condensation transition in pure states. The other perspective is to look at what remains. We find that the existence of anomalous strong $1$-form symmetries can be viewed as a guiding principle to novel mixed-state TO. We give general proof that anomalous strong 1-form symmetries imply the LRE nature of the mixed states, which manifests as deconfined anyons with nontrivial statistics. By analyzing the fusion rules and statistics of deconfined anyons, we show the possibility of realizing non-modular TO in mixed states.
As is clear from the three examples, intrinsic mixed-state TO can have or not have quantum memory, depending on whether the remaining deconfined anyons have nontrivial braiding statistics or only fermionic statistics Bao et al. (2023).
@@
We end with some open directions. Firstly, as already mentioned in Section III, it remains unclear whether $1$-form symmetries generated by fermions also guarantee bipartite long-range entanglement, or whether there exist counterexamples that are bipartite separable but have multipartite long-range entanglement Lessa et al. (2024). Secondly, a systematic classification of the intrinsic mixed-state TO is still lacking, with two main difficulties. The first is how to treat the weakly deconfined anyons with the peculiar one-way braiding statistics. The other is how to generalize the discussion of anyon statistics to mixed states away from the fixed-point models, like the one discussed in II.4. Finally, in our construction, we start from a topologically ordered state and get its descendants via noisy channels. It is also tempting to find systematic ways to prepare mixed-state topological order from short-range entangled mixed states, for example, by measuring mixed-state symmetry protected topological order Ma and Wang (2023); Lee et al. (2022b); Zhang et al. (2022); Ma et al. (2023); Lu et al. (2023).
@@ Acknowledgements.
Acknowledgments.—We are especially grateful to Ruihua Fan for many pieces of valuable advice and feedback on the manuscript. We also thank Tsung-Cheng Lu for pointing out to us the dependence of TEN on the choice of entanglement cut. We thank Yingfei Gu, Meng Cheng, Yu-An Chen, Zhen Bi, Jing-Yuan Chen, and He-Ran Wang for helpful discussions. This work is supported by the NSFC under Grant No. 12125405, National Key R&D Program of China (No. 2023YFA1406702), and the Innovation Program for Quantum Science and Technology (No. 2021ZD0302502). Z Wu acknowledges the support in part from Shuimu fellowships at Tsinghua University and the Engineering and Physical Sciences Research Council (EPSRC) fellowships.